(1 point) The matrix -1 -1 01 A = -16 0 Lk 0 has three distinct real eigenvalues if and only if -17.036 <k< 29.184
Answer #1

Given matrix, -1 -16 -1 0 07 -1 o To In order to get a distinct values, derivative of det (A -> I] is zero -1-1 -1 det [A - 1

uit 2+14+ 192 6 = 2 = 14 = or -2 } - Now, putting these values in ® and equating it with zero, we will get a values of k. (Thif x=-2 -(-29-(-2)*+166-2)+k=0 :-* = 28 30 but det [A-21] +0 k given by ② & 9 are not exact values of k. But it should lie in